OpenStudy (anonymous):

Use factoring, the quadratic formula, or identities to solve the equation. Find all solutions in the interval [0, 2π). (Round each answer to three decimal places.) cos(2x) + sin(x) = 1

OpenStudy (anonymous):

Do you recall the double angle cosine identity?

OpenStudy (anonymous):

Anything?

OpenStudy (anonymous):

Sorry, I was getting help for my homework. I believe I remember it. cos(2x)=1-sinx or something along those lines?

OpenStudy (anonymous):

I guess I should just plug and play from there if that is right...

OpenStudy (anonymous):

\[\cos(2x) = 1 - 2\sin^2(x)\] This one is nice because it puts x in terms of only sin.

OpenStudy (anonymous):

So just plug in the identity there in the equation given..

OpenStudy (anonymous):

It works out nicely. You should get 4 solutions on the interval.

OpenStudy (anonymous):

Workin' it out now. Looks a little crazy at first, so I'm trying to figure it out.

OpenStudy (anonymous):

You should end up with factors of sin(x)

OpenStudy (anonymous):

Awesome, thanks for the tip.

OpenStudy (anonymous):

No problem. Let us know what you came up with.

OpenStudy (anonymous):

It was an between interval [0, 2π): 0, π/6, 5π/6, π

OpenStudy (anonymous):

Got it right on the assignment, too. So, thanks again!

OpenStudy (anonymous):

Perfecto.